The heat required to raise the temperature of a large cup of water(for coffee ) from room temperature to boiling in approximately 100 kj. Express this quantity of heat in kilocalorues. Show your calculations.
Answer by stanbon(75887)

The heat required to raise the temperature of a large cup of water(for coffee ) from room temperature to boiling in approximately 100 kj. Express this quantity of heat in kilocalories. Show your calculations.
----
Note:: 1 joule = 0.239 calorie
------
100 kj = 100*(1000 joules) = 100(1000 (0.239 calories)) = 100*239 calories
= 23900 calories
----------
Cheers,
Stan H.
--------------